# Exploring the generalizability of the optimal 0.234 acceptance rate in random-walk Metropolis and parallel tempering algorithms

### v1

@@ 3.2 Non-Gaussian Proposal Densities
Next, we investigate the efficacy of different proposal distributions for the Metropolis algorithm. One of the conditions assumed by the 0.234 theorem of [20] is the use of a multivariate Gaussian increment proposal distribution $Q(x)=N(0,\sigma^{2}I_{d})$. The additional results in [23] depend on the assumptions of light-tailed proposals. In contrast, [12] extend this analysis by examining the behaviour of the algorithms with heavy-tailed proposals. They find that, unlike the diffusion behaviour seen with light-tailed proposals, heavy-tailed proposals lead to a different dynamic characterized by abrupt movements followed by periods of inactivity. Furthermore, [17] investigated optimal scaling results for non-Gaussian proposals; specifically, they investigated heavy-tailed proposal distributions such as the Cauchy distribution, and spherical proposal distributions of fixed radius. In their RWM simulation study with three different continuous i.i.d. product target densities, both the Gaussian and Cauchy proposals had an estimated asymptotically optimal acceptance rate of approximately 0.234 in $d=100$. They also found that a spherical proposal with a fixed radius of 2.38 achieved an optimal acceptance rate of about 0.234 in $d=2$ and $d\geq 10$.
@@
We extend this work with two different proposal distributions in $d=2,5,10,20,50$. The first proposal distribution we consider is the multivariate Laplace distribution, which may also be referred to as the double exponential distribution. We use location parameter $\mu=0$ and covariance parameter $\Sigma=\sigma^{2}I_{d}$, with $\sigma$ to be chosen. We also consider an example with the uniform proposal distribution within the closed interval $\left[x_{i}-\frac{b}{2},x_{i}+\frac{b}{2}\right]$ for each component $i$, where $b>0$ is the interval length. We use the standard multivariate Gaussian as the target density $\pi$.
Figure 3: ESJD as a function of acceptance rate for the Laplace proposal distribution, with dimensions $d=2,5,10,20,50$ (from top-left to bottom-right).
Figure 4: ESJD as a function of acceptance rate for the Uniform proposal distribution, with dimensions $d=2,5,10,20,50$ (from top-left to bottom-right).
Table 2: Experimental optimal RWM acceptance rates for Laplace and Uniform proposal distributions in different dimensions. Numerical values are the observed experimental acceptance rates maximising the ESJD.
Dimension $d$
Proposal Distribution [^1] Parameters for each proposal distribution are given in the experiment description above. 2 5 10 20 50
Laplace 0.3986 0.3097 0.2701 0.2449 0.2311
Uniform 0.3204 0.2791 0.2623 0.2435 0.2392
[…]
As observed in Figures 3, 4 and Table 2, an acceptance rate of approximately 0.234 remains efficient and maximizes ESJD even in lower dimensions ($d=10$) under both Laplace and uniform proposal distributions, and that the optimal acceptance rate tends to converge to 0.234 as the dimensionality of the problem increases. Specifically, for a multivariate i.i.d. Gaussian target distribution, an acceptance rate close to 0.234 maximizes ESJD. This finding is consistent across various proposal distributions, including the Laplace, Uniform, and Gaussian.
3.3 Non-smooth Target Densities: Continuous Hypercube
[…]
Figure 5: ESJD for various RWM acceptance rates for the continuous hypercube target distribution in dimensions $d=2,5,10,30,50,100$ (from top-left to bottom-right).
Table 3: Hypercube target experimental optimal RWM acceptance rates for different dimensions.
Dimension $d$ 2 5 10 20 30 50 100
Optimal Acceptance Rate 0.4359 0.2578 0.2095 0.1561 0.1507 0.1398 0.1331
[…]
We have a fairly similar ESJD vs acceptance rate curve in all dimensions as shown by Figure 5. Still, unlike the previous examples, which seem to be fairly consistently optimal with an acceptance rate of 0.234, we find (refer to Table 3) that the optimal acceptance rate for the hypercube target changes significantly as the dimension increases. This is because discontinuities introduce additional complexity in the target distribution, affecting the behaviour and performance of the RWM. As the dimension $d$ increases, the algorithm’s performance converges to the theoretical 0.1353 value. In this specific case, the shrinking optimal acceptance rate might be because of the unique zero or non-zero form of the target density beyond the hypercube boundary. As discussed by [18], whether a proposed move is accepted depends completely on whether the proposed state is inside the hypercube. Subsequently, when using a Gaussian proposal, the probability that at least one proposed component will lie outside the boundary increases, leading to a higher probability of instant rejection.
@@ 3.4 Multimodal Target Densities
When we have widely-separated modes, a RWM with the Gaussian proposal distribution struggles to reach the other modes in finite time regardless of the proposal scaling and the dimension of the target distribution. Therefore, for these multimodal target density experiments, we deliberately set the modes to be close enough to each other so that a RWM still mixes between the modes reasonably often.
@@ 3.4.1 The “Rough Carpet” Target Density
We explore two multimodal target densities in this section: first, we examine a simpler i.i.d. product density, and then one with inhomogeneous scaling factors. Our first multimodal distribution has the form $\pi=\prod_{i=1}^{n}f(x_{i})$, where the single component density $f$ is a one-dimensional density with three modes. We define our single-component density:
$f(x)=0.5N(x|\mu_{1},1)+0.3N(x|\mu_{2},1)+0.2N(x|\mu_{3},1)$ (1)
where $\mu_{1},\mu_{2},\mu_{3}\in\mathbb{R}$ and $N(x|\mu_{i},1)$ indicates the density of the univariate Gaussian distribution at point $x$ with mean $\mu_{i}$ and variance 1. This forms a “rough carpet”-like appearance for the target distribution when we take the product of this single-component density over many dimensions (see the visualization of the first component in Figure 7). Since we have $3^{d}$ modes, the density values at each mode may not be very high. Yet, since this is a i.i.d. product form, the theoretical result of 0.234 being the optimal acceptance rate should also apply.
Figure 7: Single-dimensional component density for the “rough carpet” target distribution.
We conduct various experiments in dimensions $d=20,30,50$ where we use the single-component density $f(x)=0.5N(x~{}|-\!5,1)+0.3N(x~{}|~{}0,1)+0.2N(x~{}|~{}5,1)$. As mentioned (refer to Figure 6), we set the modes to be reasonably close so that the standard RWM can jump between modes somewhat frequently.
Figure 8: ESJD for various RWM acceptance rates for the rough carpet product target distribution in dimensions $d=$ 20, 30, 50 (from left to right).
Table 4: Rough carpet product target distribution experimental optimal RWM acceptance rates for different dimensions.
Dimension $d$ 20 30 50
Optimal Acceptance Rate 0.2133 0.2202 0.2291
[…]
We report our results for the non-scaled target distribution (Section 3.4.1 point 1) in Figure 8 and Table 4. We find that the exact optimal acceptance rate that maximises ESJD is slightly below 0.234 in low dimensions, but that 0.234 is still quite applicable, and that the empirically optimal acceptance rate gets closer to 0.234 with increasing dimensionality. This seems consistent with [20] since our target density, despite being multimodal, is a product of a single-dimensional density in the $d$ components.
@@ 3.4.2 The “Rough Carpet” Target Density: Inhomogeneously Scaled Components
Next, we examine a multimodal distribution with a density of the form $\pi=\prod_{i=1}^{n}C_{i}f(C_{i}x_{i})$. This density is a more general form of the previous one by including inhomogeneous component-wise scaling factors $C_{i}$, where $C_{i}>0$ and $\mathbb{E}(C_{i})=1$ and $Var(C_{i})<\infty$. This inhomogeneously-scaled i.i.d. product target density was discussed by [23], which proved the extension of the 0.234 asymptotically optimal acceptance rate to this case. We use the same single-component density $f$ as the previous experiment in Section 3.4.1, independently sample the component scaling factors $C_{i}\sim\text{Uniform}[0,2]$ and report our results for the inhomogeneously scaled target distribution in Figure 9 and Table 5.
Figure 9: ESJD for various RWM acceptance rates for the inhomogenously scaled rough carpet target distribution in dimensions $d=$ 20, 30, 50 (from left to right).
Table 5: Inhomogenously scaled rough carpet product target distribution experimental optimal RWM acceptance rates for different dimensions.
Dimension $d$ 20 30 50
Optimal Acceptance Rate 0.2009 0.2174 0.2351
[…]
Yet again, we find that the optimal acceptance rate converges to about 0.234 with increasing dimensionality. While the lower-dimensional cases such as $d=20,30$ have a larger degree of deviation from the 0.234 figure than the previous homogeneously scaled case, in $d=50$ we find that the figure is very close to 0.234. A possible explanation for this is that the dimensions $d=20,30$ might be too low for the law of large numbers to apply for the inhomogeneous scaling factors; i.e. the scaling factors $C_{i}$ sampled from Uniform$[0,2]$ may not have a sample mean close enough to 1 in low dimensions, but in $d=50$ or higher the sample mean is much closer to 1.
@@ 3.4.3 A Tale of Three Mixtures
The above “rough carpet” distributions are multimodal, but notice that they still follow the general product form $\pi=\prod_{i=1}^{n}f(x_{i})$. Here, we examine a distribution that does not have this product assumption at all; instead, we have a weighted sum of densities as our target density. In this section, we examine a mixture of three Gaussians; a multimodal distribution with just three modes regardless of the number of dimensions. The general form of this density is
$\pi(x)=w_{1}N(x|\mu_{1},\Sigma_{1})+w_{2}N(x|\mu_{2},\Sigma_{2})+w_{3}N(x|\mu_{3},\Sigma_{3})$ (2)
where we have non-negative scalar weights $w_{i}\geq 0$ such that $w_{1}+w_{2}+w_{3}=1$, and $N$ is a multivariate Gaussian with mean $\mu_{i}\in\mathbb{R}^{d}$ and covariance matrices $\Sigma_{i}$.
@@
For these experiments in $d=20,30,50$, we set the weights $w_{1}=w_{2}=w_{3}=\frac{1}{3}$, the covariances $\Sigma_{1}=\Sigma_{2}=\Sigma_{3}=I_{d}$, and set the means $\mu_{1}=(\epsilon,0,0,\dots,0),\mu_{2}=(0,0,\dots,0),\mu_{3}=(-\epsilon,0,0,\dots,0)$ for some constant $\epsilon\in\mathbb{R}$, i.e. just varying the first component of the means and setting the other components of the means to 0. As mentioned (refer to Figure 6), we choose the constant $\epsilon$ in the first dimension to be some value such that jumping between modes still happens frequently enough; for the remainder of this section we use $\epsilon=5.0$. We refer the reader to Figure 10 for an example histogram and traceplot to illustrate how the Markov chain explores the three modes.
Figure 11: ESJD for various RWM acceptance rates for the three mixture distribution in dimensions $d=$ 20, 30, 50 (from left to right).
Table 6: Three mixture target distribution experimental optimal RWM acceptance rates for different dimensions.
Dimension $d$ 20 30 50
Optimal Acceptance Rate 0.2358 0.2497 0.2250
[…]
We find the ESJD vs acceptance rate trends look almost identical in dimensions $d=20,30,50$, and ESJD continues to be maximised at an acceptance rate of around 0.234 (see Figure 11 and Table 6). Clearly, even though the i.i.d. assumption does not hold, the 0.234 figure still seems to be quite applicable here.
@@ 3.4.4 A Tale of Three Mixtures: Inhomogenous Scaling Factors
We repeat this experiment with inhomogeneous scaling factors inside the covariance matrices $\Sigma_{i}$ for each mode. Concretely, each mode has covariance matrix $\Sigma_{i}=diag(C_{1},C_{2},\dots,C_{d})$ where $C_{i}$ are i.i.d. samples from the Uniform$[0,2]$ distribution such that $\mathbb{E}[C_{i}]=1$ and $\mathbb{E}[C_{i}^{2}]<\infty$. For this experiment, we use the same diagonal covariance matrix for each of the three modes. We report our results in Figure 12 and Table 7. We observe a very similar phenomenon as the inhomogeneously-scaled rough carpet density: in lower dimensions, the optimal acceptance rate is significantly lower than 0.234, but as $d$ increases, we get much closer to 0.234.
Figure 12: ESJD for various RWM acceptance rates for the three mixture distribution with inhomogeneous scaling in each component in dimensions $d=$ 20, 30, 50 (from left to right).

### v3

@@ 3.2 Non-Gaussian Proposal Densities
Next, we investigate the efficacy of different proposal distributions for the Metropolis algorithm. One of the conditions assumed by the 0.234 theorem of [29] is the use of a multivariate Gaussian increment proposal distribution $Q=\mathcal{N}(0,\sigma^{2}I_{d})$. The additional results in [33] depend on the assumptions of light-tailed proposals. In contrast, [16] extend this analysis by examining the behaviour of the algorithms with heavy-tailed proposals. They find that, unlike the diffusion behaviour seen with light-tailed proposals, heavy-tailed proposals lead to a different dynamic characterized by abrupt movements followed by periods of inactivity. Furthermore, [24] investigated optimal scaling results for heavy-tailed proposal distributions such as the Cauchy distribution, and spherical proposal distributions of fixed radius. In their RWM simulation study with three different continuous i.i.d. product target densities, both the Gaussian and Cauchy proposals had an estimated asymptotically optimal acceptance rate of approximately 0.234 in $d=100$, and a spherical proposal with a fixed radius of 2.38 achieved an optimal acceptance rate of about 0.234 in $d=2$ and $d\geq 10$. Finally, Sherlock [38] proved that proposal distributions following a geometric “shell” condition from Assumption 2.2 (6) should yield an optimal acceptance rate of approximately 0.234 when sampling from a suitable target. As both the Laplace and Uniform proposals satisfy this condition, the 0.234 rule should hold for high dimensions, and our experiments serve to verify this theoretical prediction in lower dimensions.
We experiment with two different proposal distributions in $d\in\{2,5,10,20,50,100\}$ and set the target density $\pi_{3}$ to a standard multivariate Gaussian:
$\pi_{3}(\mathbf{x})=\mathcal{N}(\mathbf{x}\mid\mathbf{0},I_{d})$ (11)
$Q_{L}(\mathbf{y}\mid\mathbf{x})\propto\prod_{i=1}^{d}\exp\left(-\frac{|x_{i}|}{\sigma}\right),\quad Q_{U}(\mathbf{y}\mid\mathbf{x})=\prod_{i=1}^{d}\frac{\mathbb{I}\left[|y_{i}-x_{i}|\leq\tfrac{b}{2}\right]}{b}$
@@
The first proposal distribution we consider is the multivariate Laplace distribution, which may also be referred to as the double exponential distribution. We use location parameter $\mu=0$ and covariance parameter $\Sigma=\sigma^{2}I_{d}$, with $\sigma$ to be chosen. We also consider an example with the uniform proposal distribution within the closed interval $\left[x_{i}-\frac{b}{2},x_{i}+\frac{b}{2}\right]$ for each component $i$, where $b>0$ is the interval length.
Table 2: Empirical optimal RWM acceptance rates across dimensions for different proposal distributions targeting a standard multivariate Gaussian.
Proposal $d=2$ $d=5$ $d=10$ $d=20$ $d=50$ $d=100$
$Q_{L}$: Laplace 0.3780 0.3036 0.2841 0.2570 0.2429 0.2377
$Q_{U}$: Uniform 0.3194 0.2516 0.2391 0.2368 0.2365 0.2316
As observed in Table 2 and the corresponding Figures 10, 11 in Appendix A, the optimal acceptance rates for both the Laplace and Uniform proposals exhibit a clear trend. In low dimensions, the optimal rate is significantly higher than 0.234, but converges towards 0.234 from above as the dimension increases. For the Uniform proposal, the optimal rate is already approximately 0.239 at $d=10$, and for both proposals, the rates are close to the high-dimensional limit for $d\geq 20$.
These empirical findings provide strong validation for the generalized theory of optimal scaling. While the foundational result of Roberts et al. [29] was derived for Gaussian proposals, the more general criteria from Sherlock [38], summarized in Assumption 2.2, extend this theory. The key requirement on the proposal is the shell condition (6) that requires the norm of the proposal vector to concentrate around a constant in high dimensions. Both the multivariate Laplace and Uniform distributions used here are spherically symmetric and satisfy this condition. The theory therefore predicts that the 0.234 acceptance rate should indeed be asymptotically optimal when using these proposals on a suitable target. Our results confirm this prediction, and characterize the finite-dimensional behaviour, demonstrating that convergence to the 0.234 optimum occurs from above for these proposals.
@@ 3.3 Multimodal Target Densities
When we have widely-separated modes, a RWM with the Gaussian proposal distribution struggles to reach the other modes in finite time regardless of the proposal scaling and the dimension of the target distribution. Therefore, for these multimodal target density experiments, we deliberately set the modes to be close enough to each other so that a RWM still mixes between the modes reasonably often. We list out the various target densities before detailing the experimental results.
@@ 3.3.1 The “Rough Carpet” Target Density
First, we examine a simpler i.i.d. product density. Our first multimodal distribution has the form
$\pi_{4}(\mathbf{x})=\prod_{i=1}^{d}f_{4}(x_{i}),\quad f_{4}(x)=0.5\,\mathcal{N}(x|\mu_{1},1)+0.3\,\mathcal{N}(x|\mu_{2},1)+0.2\,\mathcal{N}(x|\mu_{3},1)$ (12)
where the single component density $f_{3}$ is a one-dimensional density with three modes at $(\mu_{1},\mu_{2},\mu_{3})=(-5,0,5)$ and $\mathcal{N}(x|\mu_{i},1)$ indicates the density of the univariate Gaussian distribution at point $x$ with mean $\mu_{i}$ and variance 1. This forms a “rough carpet”-like appearance for the target distribution when we take the product of this single-component density over many dimensions (see the visualization of the first two components in Figure 3 for an example). Since we have $3^{d}$ modes, the density values at each mode may not be very high. Yet, since this is a i.i.d. product form, the theoretical result of 0.234 being the optimal acceptance rate should also apply.
Figure 3: Density of the first two dimensions of the “rough carpet” target distribution with MCMC samples from one RWM simulation plotted in red. One out of every 200 MCMC samples drawn are plotted.
@@ 3.3.2 The “Rough Carpet” Target Density: Inhomogeneously Scaled Components
Next, we examine a multimodal distribution with a density of the form
$\pi_{5}(\mathbf{x})=\prod_{i=1}^{d}C_{i}f_{5}(C_{i}x_{i}),~{}~{}f_{5}(x)=0.5\,\mathcal{N}(x|\mu_{1},1)+0.3\,\mathcal{N}(x|\mu_{2},1)+0.2\,\mathcal{N}(x|\mu_{3},1).$ (13)
Note that $f_{5}=f_{4}$, so this density $\pi_{5}$ is a more general form of $\pi_{4}$ by including inhomogeneous component-wise scaling factors $C_{i}$, where $C_{i}>0$ and $\mathbb{E}(C_{i})=1$ and $\text{Var}(C_{i})<\infty$. This inhomogeneously-scaled i.i.d. product target density was discussed by [33], which proved the extension of the 0.234 asymptotically optimal acceptance rate to this case. We independently sample the component scaling factors $C_{i}\sim\text{Uniform}[0.02,1.98]$.
@@ 3.3.3 A Tale of Three Mixtures
The above “rough carpet” distributions are multimodal, but notice that they still follow the general product form $\pi(\mathbf{x})=\prod_{i=1}^{d}f(x_{i})$. Here, we examine a distribution that does not have this product assumption at all; instead, we have a weighted sum of densities as our target density. We examine a mixture of three Gaussians; a multimodal distribution with just three modes regardless of the number of dimensions. The density is
$\pi_{6}(\mathbf{x})=w_{1}\mathcal{N}(\mathbf{x}|\boldsymbol{\mu}_{1},\Sigma_{1})+w_{2}\mathcal{N}(\mathbf{x}|\boldsymbol{\mu}_{2},\Sigma_{2})+w_{3}\mathcal{N}(\mathbf{x}|\boldsymbol{\mu}_{3},\Sigma_{3})$ (14)
where we have non-negative scalar weights $w_{i}\geq 0$ such that $w_{1}+w_{2}+w_{3}=1$, and $\mathcal{N}$ is the density function of a multivariate Gaussian with mean $\boldsymbol{\mu}_{i}\in\mathbb{R}^{d}$ and covariance matrix $\Sigma_{j}\in\mathbb{R}^{d\times d}$.
@@
For these experiments, we set the weights $w_{1}=w_{2}=w_{3}=\frac{1}{3}$, the covariances $\Sigma_{1}=\Sigma_{2}=\Sigma_{3}=I_{d}$, and set the means $\boldsymbol{\mu}_{1}=(\epsilon,0,0,\dots,0),\boldsymbol{\mu}_{2}=(0,0,\dots,0),\boldsymbol{\mu}_{3}=(-\epsilon,0,0,\dots,0)$ for some constant $\epsilon\in\mathbb{R}$, i.e. just varying the first component of the means and setting the other components of the means to 0. As mentioned (refer to Figure 2), we choose the constant $\epsilon=5.0$ in the first dimension after observing via traceplots that jumping between modes still happens frequently enough. We refer the reader to Figure 4 for an example histogram and traceplot to illustrate how the Markov chain explores the three modes.
@@ 3.3.4 A Tale of Three Mixtures: Inhomogenous Scaling Factors
We repeat our experiment with inhomogeneous scaling factors applied to each Gaussian mode. Concretely, we instantiate our target density as
$\pi_{7}(\mathbf{x})=\sum_{k=1}^{3}w_{k}\cdot(\prod_{j=1}^{d}C_{j})\cdot\mathcal{N}(\mathbf{x}\odot\mathbf{C}|\mu_{k},I_{d})$ (15)
where $C_{i}\sim\mathrm{Uniform}[0.2,1.8]$ and $\mathbf{x}\odot\mathbf{C}$ denotes the Hadamard (element-wise) product $(x_{1}C_{1},...,x_{d}C_{d})$. This allows each Gaussian mode to follow the functional form defined in Roberts and Rosenthal [33] Section 6; each mode has its own density of the functional form $\pi(\mathbf{x})=\prod_{i=1}^{d}C_{i}f(C_{i}x_{i})$, where $\{C_{i}\}$ are i.i.d. with $\mathbb{E}(C_{i})=1$ and $\mathbb{E}(C_{i}^{2})<\infty$.
3.3.5 Multimodal Experiments and Results
We conduct our experiments in $d\in\{20,30,50\}$ using the default experimental procedure described. We summarize our results in Table 3, and refer the reader to Appendix A Figures 12, 13, 14, and 15 for plots of the trend of ESJD with acceptance rate.
Table 3: Empirical optimal RWM acceptance rates for multimodal target distributions.
Target $d=20$ $d=30$ $d=50$
$\pi_{4}$: Rough Carpet (homogeneous) 0.2133 0.2202 0.2291
$\pi_{5}$: Rough Carpet (inhomogeneous) 0.2009 0.2174 0.2351
$\pi_{6}$: 3-Mixture (homogeneous) 0.2489 0.2454 0.2394
$\pi_{7}$: 3-Mixture (inhomogeneous) 0.2360 0.2334 0.2371
The results for our multimodal experiments provide strong empirical support for the generalizability of the 0.234 acceptance rate heuristic. We observe that for all four target distributions, the optimal acceptance rate converges towards 0.234 as the dimension increases, even for those that lack an i.i.d. product structure. Their optimal acceptance rates are already very close to 0.234 even in low dimensions ($d=10$). Theoretically, the 0.234 value is likely to be justified by the general criteria of Sherlock [38] that are stated in Assumption 2.2. The three mixture targets ($\pi_{6},\pi_{7}$), while not product densities, are constructed from isotropic Gaussians. In high dimensions, they form a well-behaved geometric shell with low eccentricity, thus satisfying the target conditions for the 0.234 rule to apply. The rough carpet targets $\pi_{4},\pi_{5}$ are i.i.d. products of a more complex multimodal density. The component-wise multimodality creates a more intricate target landscape, which may require higher dimensions before the geometric properties average out to satisfy the shell and stability conditions as effectively; this may explain why their optimal rates approach 0.234 from below.
There is an interesting note that for both families of targets, introducing inhomogeneous scaling consistently pushes the optimal acceptance rate lower in smaller dimensions. This could be explained by the Target Eccentricity condition of Sherlock [38] (mentioned in Assumption 2.2 (5)). The inhomogeneous scaling factors introduce anisotropy into the target’s curvature, creating directions with much higher curvature than others. With a standard isotropic proposal such as the multivariate Gaussian, the optimal proposal scale must shrink to avoid constant rejections in these high-curvature directions. This makes the sampler inefficient in the less-curved directions, degrading the algorithm’s overall performance. This less efficient dynamic alters the trade-off between proposal size and acceptance probability, shifting the peak of the ESJD curve to the left. Our results show that for such low-dimensional inhomogeneously scaled targets, the optimal balance is achieved at a lower acceptance rate than the 0.234 ideal. Then, as $d$ grows, the law of large numbers mitigates the effect of any single scaling factor, the target becomes effectively less eccentric, and the optimal acceptance rate recovers to 0.234.
3.4 Rosenbrock Target Densities
The multimodal experiments in Section 3.3 demonstrated the effectiveness of the 0.234 acceptance rate for multimodal distributions, where the main challenge is mode isolation. In practice, however, many posterior landscapes are not only multi-peaked but also have nonlinear correlation structure, providing challenges for local proposals even when there is only a single global mode. We examine this setting by sampling three different extensions of the two-dimensional Rosenbrock target density to $n$-dimensions [11, 9, 27], each of which have different curvature and therefore different difficulty from a MCMC perspective.
Rosenbrock [31] originally introduced the Rosenbrock function as an optimization problem: a non-convex function which has a global minimum inside a long, narrow parabolic canyon. The Rosenbrock function can then be converted into a probability density that preserves these properties. Because the local correlation structure changes rapidly along the ridge, step sizes that work near the mode are either rejected or become inefficient elsewhere. Consequently, Rosenbrock targets have become common benchmarks to test MCMC algorithms [15, 13, 14, 18]. Including these targets in our empirical study extends our findings to targets with curved correlation structure.
We first describe the unnormalized Rosenbrock target densities, then describe our experimental procedure and results. We use the Rosenbrock kernels described by Pagani et al. [27].
3.4.1 Full Rosenbrock
The $d$-dimensional Full Rosenbrock distribution [11] has the following kernel for $\mathbf{x}\in\mathbb{R}^{d}$:
$\pi_{8}(\mathbf{x})\propto\exp\left\{-\sum_{i=1}^{d-1}\left[100(x_{i+1}-x_{i}^{2})^{2}+(1-x_{i})^{2}\right]/20\right\}$ (16)
The original 2-D Rosenbrock resembles a long banana-shaped valley whose floor follows the parabola $x_{2}=x_{1}^{2}$. As stated in Pagani et al. [27], as $d$ increases, the long, narrow ridge that makes sampling challenging becomes more concentrated around the mode, decreasing the problem’s difficulty. However, the variance of $x_{d}$ increases steeply as each new random variable is directly dependent on the squared value of the previous variable.
3.4.2 Even Rosenbrock
The $d$-dimensional Even Rosenbrock distribution extended from the optimization function in [9] has the following kernel for $\mathbf{x},\boldsymbol{\mu}\in\mathbb{R}^{d}$, where $d$ must be even:
$\pi_{9}(\mathbf{x})\propto\exp\left\{-\sum_{i=1}^{d/2}\left[(x_{2i-1}-\mu_{2i-1})^{2}+100(x_{2i}-x_{2i-1}^{2})^{2}\right]/20\right\}$ (17)
The Even Rosenbrock partitions the state into $\tfrac{d}{2}$ independent 2-D Rosenbrock blocks. Because the blocks do not interact, the global target is the product of smaller bananas; the joint density still exhibits the same curved valleys within each block but lacks the long-range, serial dependence of the Full Rosenbrock. Consequently, the optimal proposal scale can adapt locally to each block, and we expect higher ESJD (and hence slightly lower optimal acceptance rates) than in the fully coupled case.
3.4.3 Hybrid Rosenbrock
The $n$-dimensional Hybrid Rosenbrock distribution introduced by Pagani et al. [27] has the following kernel for $\mathbf{x}\in\mathbb{R}^{d}$, where $\mu,x_{j,i}\in\mathbb{R};a,b_{j,i}\in\mathbb{R}^{+}(\forall j,i)$:
$\pi_{10}(\mathbf{x})\propto\exp\left\{-a(x_{1}-\mu)^{2}-\sum_{j=1}^{n_{2}}\sum_{i=2}^{n_{1}}b_{j,i}(x_{j,i}-x_{j,i-1}^{2})^{2}\right\}$ (18)
The integers $n_{1}$ and $n_{2}$ describe a grid of blocks: there are $n_{2}$ vertical stacks, each containing $n_{1}\!-\!1$ Rosenbrock-type links and a shared root coordinate $x_{1}$. The total dimension is therefore $d=(n_{1}-1)n_{2}+1$. Within every column $j$, the curvature coefficients $b_{j,i}$ can vary with depth $i$, producing steeper or flatter valleys as we move away from the root. Horizontally, however, columns are independent, so the Hybrid form interpolates between the serial dependence of the Full Rosenbrock (one long chain: $n_{2}=1$) and the independent blocks of the Even Rosenbrock ($n_{1}=3$, $n_{2}=d/2$). This allows the difficulty of the distribution to be modified by setting $n_{1},n_{2}$; larger $n_{1}$ increases non-linearity within a column, while larger $n_{2}$ raises overall dimension without adding extra curvature.
3.4.4 Simulations and Results
For our Rosenbrock simulations, we found that our initial settings of 200,000 Markov chain steps and 20 seeds were insufficient to produce meaningful trends, so we ran more simulations for more steps with the Rosenbrock targets. In this section, we continue using a Gaussian proposal with 40 different variance values. However, we run each simulation for 1,000,000 steps and we repeated each experiment with 30 different seeds instead of the usual 20. For the Full Rosenbrock, we select $d\in\{5,10,20,30\}$. For the Even Rosenbrock, we select $d\in\{4,10,20,30\}$. For the Hybrid Rosenbrock, we select $(n_{1},n_{2})\in\{(3,2),(5,2),(7,3),(8,4)\}$ which gives total dimension $d\in\{5,9,19,29\}$ respectively. We also provide a plot of the ESJD versus acceptance rate trend for the 2-dimensional Rosenbrock. All other plots of ESJD versus acceptance rate trends for these distributions are in Appendix A Figures 16, 17, and 18 for plots of the trend of ESJD with acceptance rate.